1. a.) Give an example of how to calculate a predetermined overhead rate. b.) Using the...

1.

a.) Give an example of how to calculate a predetermined overhead rate.

b.) Using the calculated POHR in Q#1 above, show how to compute the total unit cost of a product, including direct material and direct labor.

2.

Give an example of how to compute the total cost for a job using a plantwide predetermined overhead rate.

Using the example of total job cost computed above, show how to compute the unit cost of a product.
Answer #1
1.)
a.) Let Assume the Following
Estimated Overhead $ 10,000
Estimated Direct labour Hour $2,000
Predetermined Overhead rate $ 5 Per Direct Labor Hour
(10,000 / 2,000 )
b.) Let Assume the Following
Product Y per unit Data
Direct Materials $ 100
Direct Labor $ 50
Direct Labor Hour 4
Total Unit cost of Product Y $ 170 =100+50+(4*5)
2.) Let Assume the Following
Estimated Overhead $ 100,000
Estimated Direct labour Hour 25,000
Product Z Data
Units produced 1,000
Direct Materials $ 10,000
Direct Labor $ 5,000
Total Direct Labor Hours used          1,000
Predetermined Overhead rate $ 4 =100000/25000
Total Cost of Product Z
Direct Materials        10,000
Direct Labor 5,000
Applied Overhead          4,000 =1000*4
a.) Total Cost of Product Z $ 19,000
b.) Unit cost of Product Z $ 19 =19000/1000